OpenStudy (legomyego180):

Find the Taylor Polynomial:

OpenStudy (legomyego180):

Find\[P _{4}(x)\] Centered at x=0 for the function: \[f(x)=x-4\cos(x)\]

OpenStudy (legomyego180):

\[f(x)=x-4\cos(x)\] \[f'(x)=1+4\sin(x)\] \[f''(x)=4\cos(x)\] \[f ^{3}(x)=-4\sin(x)\] \[f^4(x)=-4\cos(x)\] Evaluated @ x=0 f(0)=-4 f'(0)=1 f''(0)=4 f'''(0)=0 f''''(0)=-4 \[\frac{ f^n(0) }{ n! }\] f(0)=-4 f'(0)=1/2 f''(0)=4/6 f'''(0)=0 f''''(0)=-4/24

OpenStudy (legomyego180):

I need to multiply all of these by (x-a)^n, and I think a=0 but my answer choices arent matching what Im getting

OpenStudy (legomyego180):

My answer I get is: -4+(x/2)+(2/3)x^2+0-(1/6)x^4

OpenStudy (math&ing001):

Taylor for cosin at 0 is cos(x) = 1 - x2/2! + x4/4! - x6/6!

OpenStudy (legomyego180):

Is that just something to have memorized? How do I work out this problem with that info?

OpenStudy (math&ing001):

I memorized it from using it so much, and yes you can use it, no need to work it from the beginning.

OpenStudy (legomyego180):

\[\cos(x)=1-\frac{ x^2 }{ 2! }+\frac{ x^4 }{ 4! }...\]

OpenStudy (legomyego180):

just distribute the 4?

OpenStudy (math&ing001):

Yep !

OpenStudy (math&ing001):

and add the x...

OpenStudy (legomyego180):

add the x?

OpenStudy (math&ing001):

f(x)=x−4cos(x) multiply by -4 and add x

OpenStudy (legomyego180):

So it would look like \[-4[1-\frac{ x^2 }{ 2! }+\frac{ x^4 }{ 4! }]+x\] Which gave me the answer -4 +x+2x^2-(1/6)x^4

OpenStudy (legomyego180):

Ty for your help @math&ing001

OpenStudy (math&ing001):

You're welcome @legomyego180 =) I like you username btw :P